### v1

# A Review on Single-Problem Multi-Attempt Heuristic Optimization

Abstract.
In certain real-world optimization scenarios, practitioners are not interested in solving multiple problems but rather in finding the best solution to a single, specific problem. When the computational budget is large relative to the cost of evaluating a candidate solution, multiple heuristic alternatives can be tried to solve the same given problem, each possibly with a different algorithm, parameter configuration, initialization, or stopping criterion. The sequential selection of which alternative to try next is crucial for efficiently identifying the one that provides the best possible solution across multiple attempts. Despite the relevance of this problem in practice, it has not yet been the exclusive focus of any existing review. Several sequential alternative selection strategies have been proposed in different research topics, but they have not been comprehensively and systematically unified under a common perspective.
This work presents a focused review of single-problem multi-attempt heuristic optimization. It brings together suitable strategies to this problem that have been studied separately through algorithm selection, parameter tuning, multi-start and resource allocation. These strategies are explained using a unified terminology within a common framework, which supports the development of a taxonomy for systematically organizing and classifying them.
heuristic optimization, large computational budget, algorithm selection, parameter tuning, multi-start, resource allocation, sequential decision-making.
Preprint submitted to ACT TELO July 22, 2025.
journal: TELOjournalvolume: Preprint submitted, July 22, 2025ccs: Mathematics of computing Mathematical optimizationccs: Computing methodologies Heuristic function constructionccs: Information systems Decision support systemsccs: Human-centered computing Interaction design process and methodsccs: General and reference Surveys and overviews
1. Introduction
An optimization problem consists of finding the best solution from a search space that either maximizes or minimizes a given objective function, subject to specific constraints (Arora, 2015). These problems are encountered in a wide range of disciplines, including logistics, scientific research, and engineering. In some cases, the objective is to solve multiple related problems, each defined by variations in the searh space and the objective function. Gago-Carro et al. (2024) address a location-assignment problem to find the ambulance service management that maximizes the number of emergency calls attended under different demand scenarios. In other situations, the goal is to solve a single, well-defined problem. This setting is common in real-world applications where the interest lies in optimizing a specific search space and objective function. For example, in computational biology, Grado et al. (2018) aim to identify the set of brain stimulation parameters that minimizes beta power in patients with Parkinson’s disease. In renewable energy systems, Reddy (2020) searches for the wind farm layout that maximizes annual energy production based on a given terrain and wind conditions, while Zarketa-Astigarraga et al. (2023) optimize the design of a Wells turbine to maximize its efficiency across a selected set of representative sea states. In this work, we focus on single-problem scenarios.
Heuristic and metaheuristic methods [^1] Hereafter, for simplicity, we refer to both heuristic and metaheuristic methods as heuristic methods. are commonly employed to solve such single optimization problems. In the previous examples, Grado et al. (2018) apply an adaptive Bayesian optimization algorithm, while Zarketa-Astigarraga et al. (2023) perform a genetic algorithm-based optimization. Although these problems are not always formally proven to be NP-hard, they are typically challenging to solve (Kazimipour, 2019) due to characteristics such as the high dimensionality of large-scale solution spaces (Mahdavi et al., 2015; Zhan et al., 2022; Danilova et al., 2022), non-convex and non-differentiable black-box objective functions (Kazimipour, 2019; Talbi, 2009; Zhi-Quan Luo and Wei Yu, 2006; Pardalos et al., 2021; Sala and Muller, 2020), and noise in stochastic objective functions and constraints (Fu, 2002; Amaran et al., 2016). In such contexts, heuristic methods are particularly useful, as they provide acceptable solutions within reasonable computational times, even though they do not guarantee global optimality of the returned solution.
Heuristic methods usually rely on selecting and evaluating different solutions from the search space to return the one with the best objective value. In practice, when solving a single optimization problem, the evaluation of the objective function for each solution often takes only a few seconds (Grado et al., 2018; Reddy, 2020; Zarketa-Astigarraga et al., 2023; Cutler et al., 2015; Robinson et al., 2006). With such fast evaluations relative to the available computational budget, the practitioner has a large number of objective function evaluations at its disposal, enough to execute several heuristic methods until convergence. In such cases, the intelligent use of the abundant resources available is crucial to find the best possible solution to the problem of interest, as multiple attempts with different heuristic methods are possible. Nevertheless, the sequential selection of methods becomes a challenge.
There are different strategies in the literature that address the selection of the heuristic method to be considered in each attempt to solve a single optimization problem. Some works focus on restart strategies for a single local search algorithm (López-Sánchez et al., 2014; Mathesen et al., 2021), which allows solving the same problem in multiple attempts, considering different initializations of the algorithm at each one. Luersen and Le Riche (2004) focus on the kernel density estimation-based restart of the Nelder-Mead algorithm. Auger and Hansen (2005) instead, implement a restart strategy for the Covariance Matrix Adaptation Evolution Strategy, where the initial population is resampled uniformly in the search space at each restart, and the population size is progressively increased following a predefined schedule. This allows executing each attempt not only with a newly sampled initial population, but also with a systematically varied population size parameter, which is known to significantly influence the behavior and performance of evolutionary algorithms. Other works consider the joint selection of multiple algorithms along with their configurations. Souravlias et al. (2019) address a combinatorial optimization problem by defining a strategy to distribute the available computational budget across multiple attempts, each corresponding to a different configuration of Tabu Search, Variable Neighbourhood Search, or Multistart Local Search algorithms. Finally, we identify successive halving-based strategies (Li et al., 2018; Das Dores et al., 2018), which primarily focus on defining stopping criteria for each attempt. These approaches aim to allocate more evaluations to parameter configurations of supervised learning algorithms that show greater potential, in order to converge toward the best-performing configuration that minimizes the empirical error.
The above examples show a wide diversity of strategies for addressing a single problem in multiple attempts. The diversity lies not only in the selection criteria, but also in the component that differentiates the heuristic methods selected for each attempt: a heuristic algorithm, a configuration of its parameters, an initialization or a stopping criterion. Thus, each combination of these four components defines a heuristic alternative to solve the optimization problem of interest. We have observed that the alternative selection strategies focus on solving a single problem are divided into four different topics studied separately in the literature, each one focused mainly on the selection of one component. However, all these strategies are valid for the same practical Single-Problem Multi-attempt Heuristic Optimization (SIMHO) scenario, where a practitioner faces the following question: how to sequentially select heuristic alternatives to find the best possible solution to a given optimization problem in multiple attempts?. There are reviews that give an overview of alternative selection (Kotthoff, 2016; Martí et al., 2019). However, most works focus on the selection of a single component (Locatelli and Schoen, 2016; Huang et al., 2020a), while only a few address the joint selection of two components at most (Kerschke et al., 2019; Yang and Shami, 2020; Bischl et al., 2023).
Existing reviews do not comprehensively cover strategies for practitioners facing specific SIMHO problems. To overcome this gap, this paper presents a review focused on selection strategies suitable for SIMHO, encompassing strategies that address the selection of any combination of the four components. We unify existing literature related to SIMHO within a single framework, identifying a shared structure among strategies typically studied separately and formalizing them under the same terminology. Furthermore, we define a taxonomy that characterizes all strategies within the proposed framework. To the best of our knowledge, this is the first review to unify, describe, and classify selection strategies exclusively for SIMHO while considering all four components together.
The rest of the paper is organized as follows. Section 2 classifies related alternative selection strategies, identifies in that classification those valid for SIMHO, and highlights the difference of the review proposed in this work from the perspective that already exists. Section 3 introduces the framework that sets the notation, formalizes SIMHO as an alternative selection problem, and defines the abstraction that generalizes within the same terminology the existing suitable strategies. Section 4 describes the four topics in the literature about alternative selection that address SIMHO, illustrating how all of them fit into the proposed abstraction. Section 5 defines the taxonomy and characterizes the different strategies as particular cases of the framework. Section 6 discusses the utility of the proposed framework-based review. Finally, Section 7 concludes the paper.
2. Related work
There are several selection strategies aimed at identifying the best heuristic alternative from an available set. We distinguish two groups of selection strategies based on the type of information they rely on. To address a single optimization problem, single-problem strategies use information gathered during the optimization process of that problem. In contrast, multi-problem strategies extract knowledge from a set of previously solved problems and transfer it to a new problem. In multi-problem strategies, both the problems used to extract information and the target problem usually belong to a family of related problems, where each problem is referred to as an instance.
Examples of multi-problem strategies include per-instance algorithm selection (Kerschke et al., 2019; Rice, 1976) and offline parameter tuning (Hoos, 2011; Hutter et al., 2011), in which the selection is based on a supervised model trained using features extracted from a set of problem instances and the observed performance of a set of algorithms or configurations on them. Given the features of a new instance, the model recommends the most appropriate alternative. The Irace package (López-Ibáñez et al., 2016) provides selection strategies that replace supervised models with statistical racing procedures (Maron and Moore, 1997; Birattari et al., 2002), aiming to identify algorithm configurations that perform robustly across a set of instances. Iteratively, a set of configurations is sampled from a probabilistic model and evaluated on a representative subset of instances. The best-performing configurations are selected to update the model, progressively biasing the sampling process toward more promising regions of the configuration space.
A major limitation of multi-problem strategies is that they must invest a significant number of attempts to build accurate models by solving multiple related instances prior to addressing the target problem, which often results in a time-consuming process (Eiben et al., 2007). Furthermore, their effectiveness depends on access to a sufficiently large and representative set of instances (Baudiš and Pošík, 2014; Kerschke et al., 2019), which is crucial for ensuring generalization to unseen problems. This may not be feasible in settings such as SIMHO, where only a specific optimization instance is available or the focus. Nevertheless, the multi-problem strategies remain highly valuable in contexts where users must routinely address problems belonging to the same family (Roman et al., 2016).
Hyper-heuristics (Kiraz et al., 2011; Drake et al., 2020) are high-level strategies designed to automatically generate an effective heuristic algorithm for solving a specific problem during a single run. They achieve this by chaining low-level heuristics, such as swap or shift neighborhood moves in iterative local search (Burke et al., 2010; Lima and Lobo, 2004), or crossover and mutation operators of evolutionary algorithms, as used in parameter control (Karafotias et al., 2015) and adaptive operator selection (Fialho et al., 2010a). We consider these hyper-heuristics to fall within the single-problem group, as they adapt to the particular characteristics of the target problem, considering only the information collected during its optimization process. These strategies are typically defined as an iterative process composed of three key tasks: 1) credit assignment, which scores the performance of each low-level heuristic (Fialho et al., 2010b); 2) a selection mechanism, which chooses the next heuristic to apply based on these scores; and 3) a move acceptance to determine whether the new solution should be accepted. Vermetten et al. (2023) propose a per-run trajectory-based algorithm-switching policy that relies exclusively on local landscape features observed during the optimization of the target problem. Their results demonstrate that switching between two evolutionary algorithms within a single optimization process can significantly improve performance by leveraging the strengths of each algorithm. However, there also exist multi-problem hyper-heuristics, where the selection mechanism is based on the information gathered from previously solved problems. For example, Arza et al. (2020) introduce a population-based hyper-heuristic that trains a neural network on an instance and then uses the trained model to guide the selection, and application of low-level operators across other instances of the same problem family.
The performance of hyper-heuristics mainly depends on the selection mechanism (Scoczynski et al., 2021; Bouneffouf et al., 2020; Sutton and Barto, 2018; Li et al., 2014; DaCosta et al., 2008) and the parameters on which all three tasks may depend (Doerr and Zheng, 2020a). An appropriate tuning is often necessary to obtain high-quality solutions for a specific problem (Ferreira et al., 2017; Maturana et al., 2009). For instance, Vrugt et al. (2009) define a population-based hyper-heuristic where the population size parameter is incrementally adjusted for the problem being addressed. In the SIMHO scenario, as multiple attempts are available to solve a single optimization problem, we consider single-problem hyper-heuristics as another alternative to select from for valid SIMHO strategies, as they can be used together with their parameter configuration, initialization, and stopping criterion to carry out a specific attempt.
Overall, each suitable strategy for SIMHO may focus on the selection of different components. However, existing reviews in the literature typically cover at most the selection of two out of the four components (see Table 1), as their primary aim is not to unify independently studied strategies across the different topics relevant to SIMHO. Instead, they provide broader overviews of alternative selection strategies that often include related but misaligned approaches, such as multi-problem strategies that do not apply to the SIMHO setting. For instance, Huang et al. (2020a) revise single-problem strategies for parameter configuration but also include offline parameter tuning methods, which fall outside the SIMHO scope. As a result, such reviews often lack a unified formalization and rely on fragmented taxonomies with component-specific descriptions. In contrast, this review focuses exclusively on single-problem multi-attempt strategies and introduces a common framework that formalizes the strategies using consistent terminology. This unified perspective enables a more coherent and comprehensive taxonomy, providing a clearer and more practical reference for researchers and practitioners working specifically on SIMHO, while avoiding the inclusion of strategies that are not directly applicable.
Table 1. Unification of the literature to address SIMHO.
Reviews Components
Heuristic algorithm Parameter configuration Initialization Stopping criterion
(Kotthoff, 2016) (Souravlias et al., 2021) (Kerschke et al., 2019)
(Huang et al., 2020a) (Schede et al., 2022)
(Martí et al., 2019) (Locatelli and Schoen, 2016)
(Yang and Shami, 2020) (Bischl et al., 2023) (Elshawi et al., 2019)
This work
3. SIMHO framework
An optimization problem consists of finding the best solution among a set of possible ones, which maximizes a specific performance function. Formally, it could be defined as follows
(1) $\operatorname*{\arg\max}_{x\in X}f(x)$
where $X$ is the solution space and $f$ is the objective function that measures how good a solution is [^2] Without loss of generality, we assume that the optimization problem is a maximization problem. In the case of minimization, it would be enough to replace $f(x)$ in Equation (1) with f(x)-f(x).. In this section, we define the conditions considered to solve Problem (1), and the generalization of the suitable literature strategies to address it.
3.1. Multi-attempt heuristic optimization
Certain real-world problems are best solved using heuristic optimization methods. Heuristic algorithms are optimization algorithms that provide acceptable solutions in reasonable computational times, without any guarantee of optimality. A wide variety of such algorithms have been developed over the last decades (Vermetten et al., 2024), including direct search methods (e.g., Nelder–Mead (Nelder and Mead, 1965) and Hooke-Jeeves (Hooke and Jeeves, 1961) algorithms), evolutionary algorithms (e.g., Genetic Algorithms (Glover et al., 2003), Estimation of Distribution Algorithms (Larrañaga, 2002) and population-based hyper-heuristics (Yuen et al., 2013; Fei Peng et al., 2010)) and gradient-based methods (e.g., RMSProp (Tieleman and Hinton, 2012) and ADAM (Kingma and Ba, 2014) considered when training large neural networks). With sufficiently large computational budgets, multiple executions of different heuristics can be afforded, each providing an attempt to solve the problem.
To carry out each available attempt, four components must be fixed: 1) the heuristic algorithm, 2) its parameter configuration, 3) its initialization and 4) the stopping criterion that limits its execution time. A combination of these four components represents a reproducible sequence of computational instructions (Arza et al., 2024), which we call alternative. Its execution returns the best-visited solution in terms of the objective function. The objective value of this solution determines the quality of the alternative, which can only be revealed by executing it. Formally, we define the set of alternatives as
(2) $A=\big\{(h,\theta,\xi,c)~|~h\in H,\theta\in\Theta,\xi\in\Xi,c\in C\big\}$
where $H$, $\Theta$, $\Xi$ and $C$ are the sets of possible heuristic algorithms, parameter configurations, initializations and stopping criteria, respectively (see Appendix A for a more detailed description). The contribution of each alternative $a\in A$ to the optimization is given by the information it provides after its execution, which in practice is limited to the following triplet
(3) $I_{a}=\{x_{a},t_{a},f_{a}\}$
where $x_{a}$ is its best-visited solution, $t_{a}$ is its total execution time and $f_{a}=f(x_{a})$ is its quality.
In the described context, we define the multi-attempt heuristic optimization as the search for the solution to Problem (1) with the highest possible objective value under the following conditions:
  • Availability of different alternatives. A set of alternatives with a priori unknown qualities is available [^3] In this work, we consider the set of alternatives as given and its definition is out of the scope of this paper. For common challenges and tips on defining an appropriate set of alternatives see (Souravlias et al., 2021; Kotthoff, 2016)..
  • Availability of multiple attempts. The computational budget is large enough such that different alternatives can be executed.
As the quality of each alternative is the objective value of the best-visited solution, the single-problem multi-attempt heuristic optimization can be reformulated as an alternative selection problem as follows
(4) $\operatorname*{\arg\max}_{a\in A}f_{a}$
where the alternative of the highest possible quality must be found over the multiple available executions. Thus, the reformulation shifts the exploration of the solution space $X$ to the set of alternatives $A$. The information of the same $A$ varies on different optimization problems, thus the qualities of each alternative $a\in A$ will be different for different definitions of Problem (1). Therefore, finding the best $a\in A$ to solve Problem (4) is in itself an optimization problem, where the desired solution is the best-visited solution $x_{a}$ provided by the $a\in A$ with the highest quality $f_{a}$. We define this problem generically and any approach that addresses it as SIMHO problem and strategy, respectively.
3.2. The abstract SIMHO strategy
A naive SIMHO strategy consists of iteratively executing a different alternative $a\in A$ drawn from a uniform random distribution to reveal its quality until the available executions are exhausted, and thus keeping the highest quality solution among all of them. However, this randomness does not consider the information observed in the previous executions, which may result in the selection of similar alternatives that have already shown poor quality (Mathesen et al., 2021). In the efficient SIMHO strategies identified in different topics of alternative selection, it is crucial the use of the information from previously executed alternatives to guide the selection of the remaining executions. In other words, the information observed from executed alternatives is used to define the selection preference of a new alternative.
Let $\mathcal{I}$ be the set formed by the information of executed alternatives, which we call observed information. We define the probability of an alternative $a=(h,\theta,\xi,c)\in A$ as a factorization of conditional distributions that determine the preference of each component to be selected for the next attempt conditioned on those selected before it
(5) $p(h,\theta,\xi,c)=p_{H}(h)\cdot p_{\Theta}(\theta\mid h)\cdot p_{\Xi}(\xi\mid h,\theta)\cdot p_{C}(c\mid h,\theta,\xi)$
where $p$, $p_{H}$, $p_{\Theta}$, $p_{\Xi}$ and $p_{C}$ denote the probability distributions over the set of alternatives $A$ and set of components $H$, $\Theta$, $\Xi$ and $C$, which we call distribution of alternatives and component distributions, respectively. Notice that to define the probability of $a$, its factors could use the informations from $\mathcal{I}$ associated with the executed alternatives with the same heuristic algorithm, or the same algorithm and parameter configuration, or the same algorithm, configuration and initialization (see Appendix A for a more detailed description).
Input:
  $A\colon$ Set of alternatives;
  $t_{max}\colon$ Computational budget that allows multiple attempts.
Output: $\widetilde{x}\colon$ Best solution found.
1 $\mathcal{I}\leftarrow\emptyset$;
2 $p\leftarrow$ Initialize the distribution of AA;
3 while $t<t_{max}$ do
4  $a\leftarrow$ Sample $p$;
5  $I_{a}\leftarrow$ Execute $a$ ;
6  $\mathcal{I}\leftarrow\mathcal{I}\cup\{I_{a}\}$ ;
7  $p\leftarrow$ Update probability distribution over $A$ using $\mathcal{I}$;
8
9 end while
10return $\widetilde{x}=\operatorname*{\arg\max}_{x_{a}\in\mathcal{I}}f_{a}$
Algorithm 1 Abstract SIMHO strategy
We propose an abstract definition that formalizes the existing SIMHO strategies proposed in separately studied topics under the same terminology. This consists of an iterative alternation of two processes: the information generation and the distribution update. The first focuses on expanding the observed information $\mathcal{I}$ to enable the second to update the distribution of alternatives $p$ based on it. The pseudocode associated with our proposal is shown in Algorithm 1. It starts by initializing $p$ (line 2). Then, until the available budget is exhausted, in each iteration the alternation of the two processes is applied. Firstly, to generate new information, an unexecuted alternative $a$ is sampled from $A$ based on its current distribution (line 4). Then, the selected alternative is executed to reveal its information $I_{a}$ (line 5) and add this new information to the observed information $\mathcal{I}$ (line 6). Secondly, the distribution is updated by reassigning to each unexecuted alternative a probability based on the observed information (line 7). Finally, when no more time is left to carry out a new iteration, the best-visited solution with the highest objective value among all the executed alternatives is returned (line 9) [^4] Abusing the notation, we write $x_{a}\in\mathcal{I}$ to refer solely to the third elements (best-visited solutions) of the triplets (3) that form the observed information $\mathcal{I}$. In Sections 5.3.1 and 5.3.2, the same notation is used to refer solely to the observed execution times ($t_{a}\in\mathcal{I}$) and the observed qualities ($f_{a}\in\mathcal{I}$). .
4. Topics addressing SIMHO
We have observed that the strategies suitable to address SIMHO are proposed within four different topics related to alternative selection. All of them address optimization problems in multiple attempts with a set of alternatives, and the strategies proposed in them focus on determining the alternative to be executed in each attempt until the available executions are exhausted. Specifically, each topic names the iterative process [^5] The same or similar terms are used in the literature to refer to different topics that address different objectives, as as has been shown throughout Section 2. It should be emphasized, that we chose the terms that best represent the strategies addressed by each topic, although they may be terms used for other objectives in the literature as well. depending on which of the four components of the alternative it determines. To illustrate how the proposed abstraction generalizes the existing strategies in the four topics, we describe each of them both in general terms and using illustrative examples within Algorithm 1. For this purpose, we identify in them: an optimization problem, a set of alternatives, multiple attempts and the two alternating processes.
4.1. Algorithm Selection
Algorithm Selection (AS) (Kotthoff, 2016) is motivated by the efficient allocation of computational resources between a set of heuristic algorithms that perform differently on different optimization problems. This set is also identified as Algorithm Portfolio (Souravlias et al., 2021), where it is argued that the resources are divided among its constituents. Consequently, AS iteratively specifies the heuristic algorithm in which to spend more resources.
An instantiation of our abstraction in AS topic is provided by Souravlias et al. (2019) to solve the combinatorial problem called the detection of circulant weighing matrices (Arasu and Gutman, 2010), using a portfolio consisting of two variants of the algorithms Tabu Search (Glover and Laguna, 1998), two variants of Variable Neighbourhood Search (Mladenović and Hansen, 1997), and three variants of Multistart Local Search (Lourenço et al., 2019), under different parameter configurations. In each iteration, the number of available CPUs is distributed among the portfolio constituents proportionally to their performances in previous executions, to execute again in parallel new initializations of each one for the same maximum number of evaluations. Consequently, in this case, the algorithm, parameter configuration and initialization are the components that differentiate the alternatives. Moreover, running different times multiple algorithms also involves the availability of multiple attempts. Finally, the new qualities observed when running new initializations of each algorithm and configuration pair allow updating the estimated performance for each pair and associating proportionally the number of new initializations to execute next.
4.2. Parameter Tuning
Parameter Tuning (PT) (Huang et al., 2020a) focuses on automating the parameter setting of heuristic algorithms, to facilitate non-expert users in identifying well-performing configurations and avoiding costly brute force techniques, such as grid search. This topic is closely related to evolutionary computation (Larrañaga, 2002; Glover et al., 2003), where Genetic Algorithms or Estimation of Distribution Algorithms, although originally proposed as efficient, easy to use and applicable to a wide range of problems (Holland, 2010), their effectiveness is extremely dependent on the correct selection of parameters such as population size, crossover probabilities and selection ratios (Harik and Lobo, 1999). This requires a tuning process, in which different configurations are tested to find the most appropriate one for the problem of interest. In machine learning, the topic is known as Hyperparameter Optimization (HPO) (Schede et al., 2022), where the objective is to find the hyperparameter configuration of the learning algorithm of a model, which after being learned on a training data set has the minimum loss function value on a test data set. Therefore, since it is a priori unknown which hyperparameter configuration will perform best on the training data set, the computational resources must be distributed among several learning executions carried out with different configurations. In short, PT determines in which parameter configuration to invest resources at each iteration.
Doerr and Zheng (2020a) contribute to the design of parameter-less Estimation of Distribution Algorithms within the PT topic, proposing a smart restart strategy for the compact Genetic Algorithm (cGA) (Harik et al., 1999) with exponentially increasing population size. Different population sizes of the algorithm are considered to solve a binary optimization problem. They start executing the algorithm with a small population size, which is stopped once the risk of genetic drift (Doerr and Zheng, 2020b) is deemed too high, to reinitialize the algorithm again from a uniform distribution but now associated with a larger population size. Thus, although the algorithm and stopping criterion are the same across all runs, initialization and configuration are the components defining the set of alternatives. Furthermore, the multiple restarts of the cGA with different population sizes lead to multiple attempts. Finally, in this case, the schedule to select the algorithm configuration for each execution is known in advance, since without having the information of each execution, the rule with which the population size will be updated is scheduled beforehand.
4.3. Multi-Start
Multi-Start (MS) (Martí et al., 2019; Locatelli and Schoen, 2016) is a technique used to improve the performance of heuristic algorithms, especially in non-convex problems where multiple local optima may exist. It alternates the execution of a single heuristic algorithm with its restart at another location in the solution space. Each initialization defines an independent optimization process, potentially leading to different local optima. Therefore, MS defines the new location to be considered, which determines the unknown performance of the algorithm until it is executed again.
For example, Luersen and Le Riche (2004) solve a composite laminate design problem by considering only the Nelder-Mead algorithm (Nelder and Mead, 1965) with a fixed parameter configuration. It is restarted from a different location when the stopping criterion of reaching a small, flat or degenerate simplex in one iteration of the local search is satisfied. After each run of the algorithm, knowledge of the landscape and its local optima is enriched with the best-observed solution. This is used to update a probability distribution over the solution space, which determines the probability that the algorithm from each solution can explore a new local optimum. Thus, the component that determines the set of alternatives is the initialization, running different local searches implicitly assumes the availability of multiple attempts, and the best solution from each local search generates new information that updates the probability distribution over new initializations.
4.4. Resource Allocation
Resource Allocation (RA) focuses on speeding up the evaluation of a parameter configuration or initialization of a heuristic algorithm. This is done by gradually allocating more resources to the most promising configurations or initializations and discarding the poor ones early. When different hyperparameter configurations of the same learning algorithm are evaluated, this approach is also identified as Multi-fidelity for HPO (Bischl et al., 2023; Elshawi et al., 2019; Yang and Shami, 2020). A monotonically increasing computational cost is considered for the model fitting, identifying higher fidelity evaluations with longer execution times in fits. The fidelity can be measured in terms of training epoch, number of iterations or train set size used in the previous ones for the learning algorithm. The last case carries out a Progressive sampling (Provost et al., 1999), where the sample of the entire available train set is considered and its size is progressively increased to retrain learning algorithms that have shown better results in previous trainings with smaller samples. In contrast, when evaluating different initializations of the same algorithm, the approach is related to MS and the resources are measured by the number of visited solutions. Therefore, allocating resources dynamically allows identifying unpromising alternatives quickly to have as much time as possible for those that seem to be promising according to the observed results. Thus, RA involves determining the amount of resources to be considered in each iteration.
An illustrative example of a strategy in the RA topic is the Hyperband algorithm (Li et al., 2018) designed to optimize the hyperparameters of a convolutional neural network. Hyperband simultaneously learns the weights of the neural network with different hyperparameter configurations for a few epochs at each attempt. After each batch, in the next attempt, it resumes the learning of the configurations that have led to the most promising networks for a few more epochs, and the process is repeated until only the network trained with the most promising hyperparameter configuration remains. Therefore, the optimization problem in this case consists of finding the weights of the neural network that minimize the loss function. For this purpose, different configurations of the parameters of the learning algorithm and the stopping criteria based on successive halving are considered, which define the set of alternatives. Finally, with the implicit assumption of the availability of multiple attempts, the different configurations are run with different incremental stopping criteria, to decide which configurations to continue running considering the observed performances.
5. Taxonomy of SIMHO strategies
Addressing a literature revision under the proposed framework would lead to a review that: unifies the literature more inclusively, since it considers any combination of the four components as an alternative (see the sub box on the left inside the first box of Figure 1); is specific to researchers interested in solving a single optimization problem heuristically within multiple attempts (see the middle sub box in the first box of Figure 1); and summarizes the existing literature strategies from a shared terminology and perspective, avoiding individualistic descriptions (see the sub box on the right inside the first box of Figure 1). This section focuses on completing these three outcomes. Hence, hereafter we refer to and identify any strategy in the literature that addresses SIMHO equally from our Algorithm 1.
The aim of a SIMHO strategy is to determine the distribution of attempts among the given possible alternatives. Therefore, its effectiveness mainly depends on the distribution of alternatives defined at each iteration, which is the principal difference between its possible variants. Remember that the objective of the SIMHO problem is to find the alternative with the highest quality, since this is the one that returns the best-visited solution with the highest objective value. This alternative can only be selected by the sampling, which is guided by the distribution. We have observed that in the literature, its update is carried out as follows:
  • First, it is determined how the probability distributions of the components will evolve with the observed information over the iterations. It should be considered which ones will remain unchanged, and which ones will be modified (see Rule 1 in the second box of Figure 1).
  • Then, for the distributions that are modified according to the observed information to guide the selection, the meaning that the probabilities will take to prioritize the selection is determined. A higher probability is likely to indicate a better alternative or an alternative closer to the optimum (see Rule 2 in the second box of Figure 1).
  • Finally, it is determined how to map the probabilities by making estimates from the observed information and controlling the uncertainty (see Rule 3 in the second box of Figure 1).
In the following, we define a taxonomy with three linked classification rules associated with the three previous items. Each section defines each rule in order. In them, we determine the categories considered, illustrate each category with examples, and finally classify all the reviewed works accordingly. Overall, this section provides a review that characterizes existing selection strategies abstracting from the component. In fact, we will show how the same selection strategies sometimes appear in different topics to select different components.
Figure 1. The first box (named Single-problem Multi-attempt Heuristic Optimization framework…) highlights in yellow the contributions of a review carried out throughout the proposed framework. The second box under it (named …together with a taxonomy of selection strategies) enumerates in purple the linked classification rules in blue of a SIMHO strategy, with the categories defined for each of them in green, and the contributions implied by the resulting taxonomy in yellow.
5.1. Updating criteria
Defining a SIMHO strategy to find the optimal alternative within the available large budget is challenging. By using the information observed in the previous executions to update the distribution, the sampling can be efficiently guided towards it (Mathesen et al., 2021; Li et al., 2018). This enables the efficient use of the multiple attempts available, allowing the SIMHO strategy to execute an alternative close to the highest quality. In practice, to update the distribution $p$ over $A$ using the observed information $\mathcal{I}$, it is sufficient that at least one component distribution $p_{H},p_{\Theta},p_{\Xi}$ or $p_{C}$ does so. The rest may not update, or if they do, may not require the use of the observed information for updating. In this sense, we classify the component distributions into three categories depending on the criteria used for their updating (summarized in Table 2 and illustrated in Rule 1 of Figure 1):
Fixed. The component distribution is fixed when each component value is always assigned the same probability from iteration to iteration. Thus, the distribution remains static independently of how the observed information is extended (see the first column in Rule 1 of Figure 1). In the works where the set of parameters $H$ and initializations $\Xi$ define different alternatives (Cicirello and Smith, 2005; Souravlias et al., 2019; Parsopoulos et al., 2022; Doerr and Zheng, 2020a; Auger and Hansen, 2005; El-Mihoub et al., 2022), the distribution of initializations $p_{\Xi}$ is fixed. To extend the observed information of one of the available algorithms, it is executed again considering a randomly selected initialization, i.e., an initialization sampled from a static uniform distribution defined over $X$.
Table 2. Identification of the characteristics that distinguish the three updating criteria.
Component distribution Update type Uses $\mathcal{I}$?
Static Dynamic
Fixed
Scheduled
Adaptive
Scheduled. Unlike a fixed distribution, a scheduled distribution is dynamic, although the observed information is still unused for its update. In this case, the different distributions to be considered in each iteration are predefined before executing any alternative (see the second column in Rule 1 of Figure 1). We observed this strategy for the stopping criterion distribution in some works where the set of parameters $\Theta$ or initializations $\Xi$ with stopping criteria $C$ define different alternatives. In each iteration, the same amount of resources are allocated for all non-discarded parameter configurations (Li et al., 2018; Das Dores et al., 2018; Falkner et al., 2018; Huang et al., 2020b) or initializations (Jaderberg et al., 2017), which increases exponentially or linearly from iteration to iteration, respectively. Therefore, after defining the initial amount of resources, the criterion for all iterations is completely determined, i.e., it is the initial amount raised to the power of the number of iterations or with a proportional increase to iterations. Consequently, although $C$ is formed by all possible amounts of resources, without using $\mathcal{I}$ it is known that in each iteration $p_{C}$ assigns probability 1 to the amount inferred by the number of iterations and 0 to the rest. The same applies to the configuration distribution $p_{\Theta}$ in works focused on designing parameter-free evolutionary algorithms (Auger and Hansen, 2005; Vrugt et al., 2009; Doerr and Zheng, 2020a), where increasing population size strategies are proposed.
Adaptive. As a scheduled distribution, an adaptive distribution is dynamic, but its updating is guided using the observed information. In the previous works where the set of parameters $\Theta$ and stopping criteria $C$ define different alternatives, in each iteration all the non-discarded parameter configurations are executed in parallel, i.e., it is known that all of them have been executed as many times as iterations elapsed, which allows predefining the stopping criterion distribution. However, there are also works that require the observed information to update this distribution. Huang et al. (2019) only execute one of the non-discarded configurations per iteration, so they consider individual geometric scheduling (Provost et al., 1999) for each of them. In this case, as the selected configuration for the current iteration has not been executed in all previous iterations, the number of times it has been executed can only be obtained from the history of configurations executed in each previous iteration, thus using $\mathcal{I}$. Klein et al. (2017) instead, rather than conditioning $p_{C}$ by the set of discrete $\Theta$ parameter configurations, they treat this set continuously and define for all configurations the same discrete set of possible stopping criteria $C$ formed by different train set sizes. Finally, a Bayesian optimizer is applied that uses the observed information to determine at each iteration the configuration together with the most promising stopping criterion.
5.2. Quality meaning of probability in adaptive criterion
The adaptive updating criterion is the only one that considers the observed information (see Table 2). Thus, unlike the fixed and scheduled criteria, the adaptive one allows to guide the sampling towards the best quality alternative efficiently. We have defined the probability of an unexecuted alternative as the preference with which the alternative will be sampled against the others. In practice, the adaptive criteria give to these preferences a meaning associated with the quality of the alternatives. Specifically, we distinguish two different meanings of the probabilities related to quality (see Rule 2 of Figure 1):
Table 3. Classification of the papers depending on the criterion used to update each component distribution and the quality meaning of probability derived from the adaptive ones.
Paper Updating criterion Quality meaning of $p$
$p_{H}$ $p_{\Theta}$ $p_{\Xi}$ $p_{C}$ EMQ EQI
(Mathesen et al., 2021) (Luersen and Le Riche, 2004) (Voglis and Lagaris, 2009) (Meng et al., 2023) (Wang et al., 2023) - - Adaptive -
(Addis et al., 2005) (Siddiqui et al., 2025) - - Adaptive -
(Auger and Hansen, 2005) (Doerr and Zheng, 2020a) - Scheduled Fixed -
(El-Mihoub et al., 2022) Fixed - Fixed -
(Vrugt et al., 2009) - Scheduled Adaptive -
(György and Kocsis, 2011) - - Adaptive Adaptive
(Cicirello and Smith, 2005) Adaptive - Fixed -
(Souravlias et al., 2019) Adaptive Adaptive Fixed -
(Parsopoulos et al., 2022) Fixed - Fixed Adaptive
(Jaderberg et al., 2017) - Fixed Adaptive Scheduled
(Li et al., 2018) - Adaptive - Scheduled
(Falkner et al., 2018) (Huang et al., 2020b) - Adaptive - Scheduled
(Huang et al., 2019) - Adaptive - Adaptive
(Klein et al., 2017) - Adaptive - Adaptive
(Schmidt et al., 2020) Adaptive Adaptive - -
(Thornton et al., 2013) Adaptive Adaptive - -
(Li et al., 2020) Adaptive Adaptive - -
(Das Dores et al., 2018) Adaptive Adaptive - Scheduled
(Hu et al., 2022) Adaptive Fixed Adaptive - -
Expected maximum quality (EMQ). In some papers, higher probabilities are associated with the alternatives expected to have the highest quality considering the observed information. Therefore, in each iteration, the alternative with the highest predicted quality is preferred. For example, Souravlias et al. (2019) assign more CPUs to the pairs of algorithms and parameter configurations with higher average qualities observed in their previous executions with different initializations. Thus, the probability of the alternatives with the pair of algorithm and configuration of higher estimated quality is greater.
Expected quality improvement (EQI). In other papers instead, higher probabilities are assigned to alternatives that are expected to improve on the highest quality observed so far among the executed alternatives. Consequently, the improvement in quality is followed to reach the maximum. A clear example is given by Luersen and Le Riche (2004). They assign higher probabilities to initializations that are believed to yield the exploration of new local optima not visited by the Nelder-Mead algorithm in its previous executions. In this context, exploring a region of the solution space associated with a different local optimum is desirable, as it is more likely to improve the best-observed solution.
We have only referenced some papers throughout the previous Section 5.1 and the current section to illustrate the proposed categories. However, in Table 3, we classify all the papers considered in this review according to these categories.
5.3. Two-stage process for adaptive criterion
The dependence of the adaptive criterion on the observed information requires knowing how to use it to define the update of the distribution iteratively. In the literature, we identify a two-stage process for the updating of an adaptive component distribution based on the observed information (see Rule 3 of Figure 1):
Stage 1: Prediction of utility. Firstly, to assign a probability to each value of the component, its utility is computed, i.e., the observed information is used to predict how useful the knowledge about the information of a new alternative with that component value would be in the search for the highest quality alternative. Depending on which of the three variables forming the observed information $\mathcal{I}$ are used, i.e., the best-visited solutions $x_{a}$, the execution times $t_{a}$ or the qualities $f_{a}$ of executed alternatives $a\in A$, the utility prediction takes different meanings that are described in the following two sections (summarized in the Table 5).
Stage 2: Control of uncertainty. Secondly, the uncertainty of the predicted utilities must be controlled to determine the final component distribution. When the number of executed alternatives is low, the representation of the total information is poor, leading to uncertain utility predictions. In this case, the probabilities should be less biased by the utilities to avoid ignoring relatively different alternatives with promising information. Consequently, as the amount of the observed information increases, the representation of the total information improves, reducing the uncertainty of the utilities (see the third column in Rule 1 together with Rule 3 of Figure 1). Therefore, the uncertainty should be controlled considering the resources consumed so far by the SIMHO strategy, i.e., the execution time elapsed or the number of iterations spent in obtaining the observed information. This can be carried out in an explicit or implicit manner, depending on whether the resources consumed interact analytically with the utility in defining the probability or not, respectively. In the remainder of this section, we describe the different classifications that are given in practice (summarized in Table 6).
The following sections exemplify the two-stage process depending on the meaning to be considered for the probability. In each section, we describe how the utility is predicted and how its uncertainty is controlled in practice (possible settings of Table 4).
Table 4. Characteristics to be considered for the update of an adaptive component distribution.
Adaptive component distribution
How to predict utility? How to control uncertainty?
Determining the use of the observed information $\mathcal{I}=\big\{\hskip-5.69054pt\underbrace{(x_{a},t_{a},f_{a})}_{\textstyle\textit{Which to use?}}\hskip-5.69054pt~|~a\text{ is executed}\big\}$ Considering Explicit/ Implicit interaction of consumed resources with utility to define probability
5.3.1. Probability related to expected maximum quality
One of the options for guiding sampling toward the highest-quality alternative is to assign higher probabilities to the component values expected to define the highest-quality alternatives. In the papers considering this association of quality with a component distribution, we observed that the utility predictors are based only on the qualities of the observed information (the $f_{a}\in\mathcal{I}$) or these are combined with their execution times (the $t_{a}\in\mathcal{I}$). Moreover, the uncertainty of the predicted utilities is measured either explicitly or implicitly.
  • Utility prediction with $f_{a}$. Among the works that only use the qualities, we find those that estimate the qualities in a pointwise manner, calculating weighted averages of the $f_{a}$ observed in the previous executions of each algorithm (Souravlias et al., 2019; Parsopoulos et al., 2022), or using their maxima as an estimate of the quality of a parameter configuration (Li et al., 2018; Falkner et al., 2018; Huang et al., 2020b) or an initialization (György and Kocsis, 2011; Vrugt et al., 2009; Siddiqui et al., 2025). Other works instead, carry out density estimations. Cicirello and Smith (2005) first estimate the Gumbel distribution of the maximum quality of each algorithm, and then calculate the expected value using that distribution. Addis et al. (2005) determine the restart position for the local search using a smooth approximation of the objective function. They define a local Gaussian kernel estimate of the piecewise constant function represented by the local optima and its objective values. Since only the observed local optima from the executed initializations are known, the utility of a new initialization is approximated using the kernel estimate restricted to the qualities of the nearby explored starting points.
  • Utility prediction with $f_{a}$ and $t_{a}$. Regarding the works that also use the execution times, we still find point estimators that use the times in addition to the qualities to calculate averages (Hu et al., 2022). However, we also find interval estimates. They consist of defining an upper and lower bound of the quality of algorithms (Li et al., 2020) or parameter configurations (Huang et al., 2019) using the $f_{a}$ and $t_{a}$ observed for each of them. Moreover, we found works with more elaborate proposals, which try to model the learning curve (Schmidt et al., 2020) or convergence curve (Hu et al., 2022) of the hyperparameter optimizer of an algorithm, to consider its extrapolation in the remaining execution time as the utility prediction of each algorithm.
  • Explicit uncertainty control. Among the techniques with explicit control of uncertainty, we find Boltzmann exploration (Cicirello and Smith, 2005) and decay $\varepsilon$-greedy (Schmidt et al., 2020). Their respective temperature parameter and $\varepsilon$ decrease with the time or iterations elapsed in all previous executions of the algorithms, allowing the probabilities assigned to each algorithm to get closer to their predicted utilities over iterations. Moreover, methods based on upper confidence bounds (UCB) explicitly add to each utility an amount inversely proportional to the time or iterations spent in each algorithm (Schmidt et al., 2020; Hu et al., 2022) or initialization (György and Kocsis, 2011). Therefore, higher utilities of algorithms or initializations in which more resources are invested are equalled to smaller predicted utilities for those with less investment.
  • Implicit uncertainty control. For implicit control of uncertainty instead, we find techniques that are based on discarding. At each iteration the algorithms, parameter configurations or initializations that have a significantly worse utility than the rest are discarded, e.g., an estimated quality interval with an upper bound below the lower bound of the rest (Li et al., 2020) or a maximum quality value below the top percentage in the ranking (Li et al., 2018; Das Dores et al., 2018; Jaderberg et al., 2017; Falkner et al., 2018; Huang et al., 2020b). Notice that in these cases, the number of algorithms, configurations or initializations not discarded that will be executed again in the next iteration decreases with the growth of iterations, increasingly allocating more resources of an iteration to those with higher predicted utilities. In addition, Huang et al. (2019), after discarding configurations by comparing their estimated confidence intervals of quality, compare the interval shortening rates of the rest to select a unique configuration to be executed in the next iteration.
Table 5. Classification of papers depending on the meaning of the utility predictor and the observed information used for it.
Paper Utility predictor Which to use from $\mathcal{I}$?
$x_{a}$ $t_{a}$ $f_{a}$
(Souravlias et al., 2019) (Li et al., 2018; Das Dores et al., 2018) (György and Kocsis, 2011) (Parsopoulos et al., 2022) (Jaderberg et al., 2017) (Falkner et al., 2018) (Huang et al., 2020b) (Vrugt et al., 2009) (Siddiqui et al., 2025) Point estimation
(Hu et al., 2022)
(Huang et al., 2019) (Li et al., 2020) Interval estimation
(Mathesen et al., 2021) (Meng et al., 2023) (Cicirello and Smith, 2005) (Klein et al., 2017) (Thornton et al., 2013) (Addis et al., 2005) (Falkner et al., 2018) (Huang et al., 2020b) Density estimation
(Schmidt et al., 2020) (Hu et al., 2022) Extrapolation
(Luersen and Le Riche, 2004) (Voglis and Lagaris, 2009) (Wang et al., 2023) Exploratory probability
  • In contrast, Souravlias et al. (2019) allocate the number of available CPUs among the different algorithms proportionally to their pointwise estimates of quality. Parsopoulos et al. (2022) instead, dedicate each worker node to a portfolio algorithm and use an adaptive pursuit strategy to allocate probabilistically the resources invested in each parallel batch based on the point estimates of quality of each algorithm in previous batches. Therefore, without discarding any algorithm, the number of iterations spent in all of them regulates the uncertainty of the utilities, i.e., the estimates become more certain with iterations, allocating more resources to the most promising algorithms. Moreover, Schmidt et al. (2020) consider the $\varepsilon$-greedy technique in which $\varepsilon$ is the invariant preference defined at all iterations for all algorithms that are not the most useful. Consequently, the uncertainty of their utilities is measured in a constant manner, without discarding any algorithm for having assigned as minimum preference $\varepsilon$.
    Other works related to evolutionary computation instead, consider random sampling techniques to control the uncertainty associated with the initialization of the population. They randomly initialize population-based algorithms resulting from hyper-heuristics (Vrugt et al., 2009) or quantum-inspired evolutionary algorithms (Siddiqui et al., 2025), while incorporating a subset of the previously visited best solutions. These strategies aim to strike a balance between retaining promising solutions and exploring new regions of the search space. Thus, uncertainty is controlled by that portion of the randomly initiated population.
Table 6. Classification of the papers depending on how they control the uncertainty of utilities.
Paper Uncertainty controller How is the control?
Explicit Implicit
(Cicirello and Smith, 2005) Boltzmann exploration
(Schmidt et al., 2020) Decay $\varepsilon$-greedy
(Schmidt et al., 2020) (Hu et al., 2022) (György and Kocsis, 2011) UCB + argmax
(Schmidt et al., 2020) $\varepsilon$-greedy
(Huang et al., 2019) (Li et al., 2020) Lower-upper bound-based
(Li et al., 2018) (Das Dores et al., 2018) (Falkner et al., 2018) (Huang et al., 2020b) Successive halving
(Souravlias et al., 2019) CPU allocation
(Parsopoulos et al., 2022) Adaptive pursuit resource allocation
(Mathesen et al., 2021) (Meng et al., 2023) (Klein et al., 2017) (Thornton et al., 2013) (Falkner et al., 2018) (Huang et al., 2020b) Acquisition function + argmax
(Luersen and Le Riche, 2004) (Jaderberg et al., 2017) (Voglis and Lagaris, 2009) (Addis et al., 2005) (Vrugt et al., 2009) (Wang et al., 2023) (Siddiqui et al., 2025) Random sampling
5.3.2. Probability related to expected quality improvement
The second option for guiding sampling toward the optimal alternative consists of associating higher probabilities to the component values expected to define alternatives with more chances for improving the best observed quality. For a component distribution that defines probabilities with this meaning, we found works that continue defining the utility predictor considering the qualities of the observed information (the $f_{a}\in\mathcal{I}$), although we also observed the use of only the best-visited solutions to predict utilities (the $x_{a}\in\mathcal{I}$). However, in all of them, it is always considered an implicit control of uncertainties.
  • Utility prediction with $f_{a}$. The literature techniques that predict the utilities using the qualities correspond to the first stage of Bayesian optimization, i.e., the prediction by a model. Utility predictors are models that capture the dependence between the component value and the $f_{a}$ of the alternatives they define. The Gaussian process is the most commonly used surrogate model, which estimates a distribution for the quality of each unexecuted parameter configuration (Klein et al., 2017; Hu et al., 2022) or initialization (Mathesen et al., 2021; Meng et al., 2023) or stopping criterion (Klein et al., 2017), after training the model with the pairs formed by the executed component values and their observed qualities. Thornton et al. (2013) instead, use random forest models to predict simultaneously the quality of algorithm and parameter configuration combinations, arguing that they are good predictors with discrete and high-dimension input data. In other works (Falkner et al., 2018; Huang et al., 2020b), an adaptation of the Tree Parzen Estimator (Bergstra et al., 2011) is used to model densities over the executed parameter configuration space.
  • Utility prediction with $x_{a}$. As an example of utility prediction that uses the best-visited solutions, Luersen and Le Riche (2004) estimate the density of the set formed by the pairs of starting points $\xi$ and local convergence points $x_{a}$ visited in each previous execution of the Nelder-Mead algorithm. They argue that the topology of the basins of attraction can be efficiently summarized by using only these two points among all the points visited during the local searches. The utility predictor is the density estimator obtained by applying a Gaussian Parzen-Window. Thus, the predicted utility for each new initialization is the probability that the Nelder-Mead departed from that point can explore a different region of the solution space and converge to a new local optimum. Similarly, Voglis and Lagaris (2009) estimate the probability that an initial point does not belong to the region of attraction of any of the previously identified minima. These probabilities depend on the distances between the initial points and the local minima found by the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm (Rinnooy Kan and Timmer, 1987), the direction of the gradient, and the frequency with which each minimum has been discovered. Wang et al. (2023) consider a similar potential metric for evolutionary computation methods. They show how the population of evolutionary methods with simple restarts tend to repeat previous evolution processes ending in the basins of attraction already explored. To address this issue, they use the evaluated solutions to divide the search space into hill regions for estimating the basins of attractions, and encourage further exploration in those of higher potential for finding unidentified ones.
  • Implicit uncertainty control. Although we only found techniques that control the uncertainty of the utilities implicitly, we distinguish some that consider acquisition functions and others that apply random samplings.
    When using a Bayesian optimization, the uncertainty of the quality distributions predicted by the model for each component value is controlled by an acquisition function. Among the most popular options is the expected improvement (Mathesen et al., 2021; Meng et al., 2023; Falkner et al., 2018; Huang et al., 2020b). It maps to each component value the expected quality increment over the best observed, which is larger for values predicted with higher mean or variance of quality. Klein et al. (2017) also use the more recent entropy search acquisition function. It favours component values with larger expected information gains about the optimum, rather than trying to evaluate close to the optimum. Notice that the larger the observed information, the predicted distributions have a lower variance, leading to closer acquisition functions to the predicted model. Thus, the number of iterations implicitly measures the uncertainty.
    Other works instead, use a random sampling technique to control the uncertainty of the predicted exploratory probabilities. Luersen and Le Riche (2004) randomly select a subset of unexecuted initializations, and then choose the one with the highest predicted probability among them to reinitialize the algorithm from that location. Voglis and Lagaris (2009) uniformly sample starting points until they find one with a probability of not belonging to a region of attraction that is greater than a randomly selected number from a uniform distribution between 0 and 1. The first starting point that meets this condition is used for the next execution of the BFGS algorithm. Therefore, these techniques allow selecting initializations that are not the best according to the predicted utilities, although it is still guided by them. Moreover, as the considered samplings are always random, the uncertainty of the predicted probabilities is measured constantly.
6. Framework utility and design of new strategies
In the previous section, we characterized the valid strategies for SIMHO within the proposed framework. This review is summarized in Table 3, Table 5 and Table 6. The combined use of these three tables enables practitioners to identify, choose and design appropriate SIMHO strategies by navigating through the components of the framework. Given the flexibility of the proposed framework, this review can be naturally extended by incorporating in the summary tables other selection strategies related to the literature topics of algorithm selection, parameter tuning, multi-start, and resource allocation.
The tables provide easy access to selection strategies associated with different components. For example, a reader interested in multi-start (MS) strategies can identify relevant works in Table 3 by checking which rows are filled in the $p_{\Xi}$ column. The corresponding section for each strategy can then be located using the last columns of Table 3, Table 5, and Table 6, which indicate where each strategy is discussed. Based on this information, the reader can gain a general understanding of the main characteristics of each strategy and choose the one that best aligns with their specific requirements and preferences. Then, the reader may identify several MS strategies suitable for their problem, such as randomized restarts (Cicirello and Smith, 2005), Bayesian optimization-based restarts (Mathesen et al., 2021), or kernel density estimation-based strategies (Luersen and Le Riche, 2004), and gain a detailed understanding of each approach by consulting the corresponding references in the literature. Moreover, since an alternative is defined as any combination of the four components, some readers may also be interested in strategies that address the selection of multiple components simultaneously. For instance, a reader seeking both parameter tuning (PT) and multi-start (MS) may combine a configuration selection strategy from one work with an initialization selection strategy from another. Alternatively, the reader may adapt strategies originally developed for one component to reuse them in the selection of a different component of interest. Indeed, we observe that the same strategies are often applied to different components across various works. For example, Bayesian optimization is employed for both configuration (Klein et al., 2017; Hu et al., 2022) and initialization (Mathesen et al., 2021; Meng et al., 2023), while exponential-increasing strategies are associated with both configuration (Auger and Hansen, 2005; Doerr and Zheng, 2020a) and stopping criteria (Huang et al., 2020b; Li et al., 2018).
The last two utilities described for the review, namely combination and reuse, can lead to the design of new strategies, which would automatically be classified within our framework. Therefore, any newly designed strategy that can be characterized by our taxonomy is considered a valid strategy for addressing SIMHO. As a result, the proposed taxonomy facilitates both the development and identification of suitable strategies, including all new or existing approaches that fall within its structure. Furthermore, the Updating criterion column in Table 3 allows us to identify underexplored types of strategies and guide future research to address these gaps. Our review shows that 84% of the included papers propose strategies for selecting one or two components, whereas only 16% address the selection of three components. This indicates that new and improved strategies could be those that consider the selection of more components, or even all four, simultaneously, as shown in recent experimental analyses (El-Mihoub et al., 2022).
7. Conclusion
This paper presents a focused review of SIMHO, which refers to the scenario of single-problem multi-attempt heuristic optimization. In this scenario, a practitioner seeks to find the best possible solution to a specific optimization problem, relying on heuristic methods and using only the information gathered throughout its optimization processes. Heuristic methods are often preferred in practice due to the challenging nature of certain optimization problems, which frequently involve objective functions that are large in scale, non-convex, or affected by noise. The performance of heuristic methods depends on four key components: the algorithm, its parameter configuration, the initialization, and the stopping criterion. The execution of a different combination of these components, known as an alternative, leads to a different optimization process of the same given optimization problem. When the available computational budget is sufficiently large, multiple attempts are allowed, making it possible to execute a different alternative in each one. In such cases, the sequential selection of alternatives plays a critical role in efficiently identifying the best combination of the four components and, ultimately, in finding a high-quality solution to the optimization problem.
Although SIMHO is commonly encountered in practice, to the best of our knowledge, this is the first review that focuses exclusively on this problem and unifies alternative selection strategies that address the selection across all four components. Selection strategies suitable for SIMHO have been studied in isolation within the topics of algorithm selection, parameter tuning, multi-start, and resource allocation. Existing reviews fail to provide a comprehensive view, either because they do not consider strategies from all four topics or because they include others that are not suitable for SIMHO scenario. As a result, practitioners aiming to address SIMHO face limitations in understanding and applying the broad range of existing strategies. This review addresses that gap by bringing together strategies that, although developed independently, are all suitable for tackling SIMHO. We formalize these strategies within a unified framework using consistent terminology, and introduce a taxonomy that organizes and characterizes them accordingly. This framework-based review thus facilitates both the identification and the development of selection strategies for SIMHO.
The framework proposed in this review unifies and conceptually describes the existing alternative selection strategies as equally valid for SIMHO. This conceptual framework could serve as the foundation for a dedicated software library designed to address SIMHO. Such a library would enable the integration and combination of diverse selection strategies from the literature for targeting different components, and maybe would include options for comparative performance analysis on different problems. This tool could fulfill a role similar to that of HyFlex (Ochoa et al., 2012), which has supported the development of improved hyper-heuristics (Alanazi, 2016; Drake et al., 2020; Ferreira et al., 2017, 2015; Sabar et al., 2015).
Appendix A Hierarchical dependence between components
In practice, there is a hierarchical dependency between heuristic algorithms and their configurations, initializations and stopping criteria. For example, two different algorithms, such as the Nelder-Mead (Nelder and Mead, 1965) direct search algorithm or the evolutionary Genetic Algorithms (Glover et al., 2003), do not share the same spaces of the last three components. In the first one, the reflection, expansion, contraction and reduction coefficients of the simplex are part of the configuration, the initialization is a single solution of the solution space and the stopping criteria are related to the shape and size of a simplex. For the second one, these components do not apply. Among the configuration are the selection, crossover and mutation operators of a population, the initialization is a set of solutions that forms the initial population, and the stopping criteria can be associated with the statistics in the population. Therefore, to be exact, we should denote $\Theta_{h}$, $\Xi_{h}$ and $C_{h}$, but for simplicity, we omit the subindex by implicitly assuming that $\Theta=\cup_{h\in H}\Theta_{h}$ and so on.
This dependence is also implicit in Definition (5) of the probability distribution of alternatives. We express it as a chain of conditional distributions, where the choice of each component is conditioned on those selected before it. For each $h\in H$ the support of the component distributions $p_{\Theta}$, $p_{\Xi}$ and $p_{C}$ is the subset $\Theta_{h}$, $\Xi_{h}$ and $C_{h}$, respectively. For example, it is always satisfied that $p_{\Theta}(\theta\mid h)=0$ for all $\theta\notin\Theta_{h}$, while $\sum_{\theta\in\Theta_{h}}p_{\Theta}(\theta\mid h)=1$.